OpenStudy (anonymous):

topic; function,domain,range if f(x) = 3 * find a. f(2x-1)/f(2x+1) b. f(2x-1)-f(2x+1)

OpenStudy (anonymous):

\[if f(x) = 3 ^{x} find a. f(2x-1)/f(2x+1) b. f(2x-1)-f(2x+1\]

OpenStudy (campbell_st):

ok... so substitute x = 2x -1 then x = 2x + 1 so you have \[\frac{f(2x -1)}{f(2x + 1} = \frac{ 3^{2x - 1}}{3^{2x + 1}}\] now apply the index law for division \[3^{(2x - 1) - (2x + 1)} = 3^{ 2x - 2x - 1 -1} = \] now simplify

OpenStudy (anonymous):

whats index law